OpenStudy (anonymous):

Why is the limit 4 and not "does not exist"?

OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x approaches two of quantity x squared minus four divided by quantity x minus two.

OpenStudy (anonymous):

\[\lim \frac{ x^2-4 }{ x-2}\]

OpenStudy (anonymous):

\[\lim_{x\to 2}\frac{x^2-4}{x-2}\]

OpenStudy (anonymous):

yes that's right

OpenStudy (anonymous):

factor, cancel, plug in two for \(x\)

OpenStudy (anonymous):

ohhh right i forgot to expand the top

OpenStudy (anonymous):

let me see if i get it give me a sec

OpenStudy (anonymous):

it is not "does not exist" because... it does exist the top is already expanded, you need to factor it

OpenStudy (anonymous):

I got 4 thank you for letting me know what I did wrong

OpenStudy (anonymous):

yw